Project

General

Profile

Kaon LT Meetings » mtg_22apr06.txt

Garth Huber, 04/07/2022 02:46 PM

 
1
		Apr 06/22 PionLT/KaonLT Analysis Meeting Notes
2
                ----------------------------------------------
3
                  (thanks to Stephen for providing 1st draft)
4

    
5
Participants:
6
 - Richard Trotta, CUA
7
 - Pete Markowitz, FIU
8
 - Dave Gaskell, JLab
9
 - Julie Roche, Ohio
10
 - Nathan Heinrich, Garth Huber, Muhammad Junaid, Stephen Kay, Vijay Kumar,
11
   Love Preet, Ali Usman, Regina
12

    
13
Please remember to post your slides at:
14
https://redmine.jlab.org/projects/kltexp/wiki/Kaon_LT_Meetings
15

    
16
Richard
17
-------
18
- 10.6 GeV Heep Singles analysis and SIMC comparison
19
  - Yield equation defined in notes with each term clarified
20
  - Naming scheme issue
21
  - Running HMS singles settings for 10.6 GeV
22
  - Shifts in W distributions from SIMC for THshms=13.80,16.25deg, but
23
    THshms=18.00,19.85 have no shift
24
    - also, the THshms=18.00,19.85deg settings have good normalization
25
      agreement with SIMC, so very encouraging
26
    - need to investigate more carefully the causes of the TH=13.80,16.25
27
      discrepancies
28
- *Suggestions*
29
  - Need to zoom in plots a bit, remove cut at 1 GeV on W, and show a line at
30
    the cut value instead
31
  - Need to add peak integrals
32
  - Pmiss components for HeeP singles?
33
    - Not as good as Coin info, but could point in the right direction
34

    
35
- *Plan for next meeting*
36
  - HMS Heep Singles at 6.2, 8.4 GeV
37
  - Then determine Offsets
38
  - Analysis breakdown of work so far
39

    
40
Ali
41
---
42
- Re-organisation of PionLT/KaonLT repos
43
- database file changes
44
  - Point to param files in new structure
45
- standard.kinematics fixes
46
  - Angles, momenta for HeeP Coin settings (6.2, 8.4 and 10.6)
47
  - Three different beam energy values for 10.6 GeV settings!
48
    - None of them were right!
49
    - All adjusted to 1 MeV for beam energies
50
    - *Should cross-check beam energies with Mark's webpage on wiki*
51
    - Angles to 3 decimal points of degree
52
- UTIL_BATCH
53
    - New scripts to run KaonLT analysis scripts
54

    
55
- *Plan for next meeting*
56
  - HeeP Singles kinematics checking followed by Physics settings
57
    - will coordinate with Richard (soon?!)
58
    - ~Next meeting should be mostly done
59
    - End of April/Early May probably reasonable
60

    
61
Vijay
62
-----
63
- Physics analysis and SIMC comparison for 4.9 GeV physics setting, all three
64
  THshms settings
65
  - Slight MMK offset
66
    - Energy has been corrected in std.kin, but not the angle
67
  - SIMC yield multiplied by factor of 7
68
  - Central regions of plots look fairly well behaved, but edges are off at
69
    times
70
  - EMiss looks a bit off on the high end
71
    - Probably due to difference in delta? SHMS and HMS have an odd discrepancy
72
      in the low delta region, which is probably due to the t-dependence in the
73
      model being off compared to data
74

    
75
- *Plan for next meeting*
76
  - Will look at low Q2 PionLT data next
77

    
78
Jacob 
79
-----
80
- Met Doug Higinbotham last week, expect new beam schedule this week (Friday)
81
  - We will get less beam time than last estimate, but higher current
82
  - Should expect 80uA
83

    
84
- SHMS Hodoscope calibrations
85
  - Work from Carlos/Dave/Jacob uploaded to hallc_replay Github
86
  - Batch script issue
87
  - Calibration is significantly improved from earlier version
88
    - previously, we were seeing bands at beta~0.5
89
    - Beta more centred at 1 now
90
    - Dave M adjusted a hard cut on P.hod.1y.GoodPosTdcTimeWalkCorr
91
      - earlier cut was too tight, eliminating some hodoscopes from the
92
        calibration
93
      - Dave G: Should the hard coded cut be a parameter? Will it need to be
94
        changed again in future?
95
        - not sure, the hope is that if it is sufficiently wide, it won't need
96
          further adjustment
97
    - length of calibration script is doubled from before, not sure exactly
98
      what else was changed
99

    
100
    - Time walk correction implemented, improved
101
      - TDC theshold parameter changed from 120 to 1200, 12000 was also tried 
102
      - Nathan: this variable is only used to make the time walk plots and has
103
        no bearing on the actual applied threshold
104
      - changes the time walk fit values, but not necessarily the goodness of
105
        fit
106

    
107
Junaid
108
------
109
- JLab account issues
110
  - No longer an active user?
111
  - Virtual user only allowed to register for 1 year?
112

    
113
- Working on high epsilon run plan
114
  - HeeP settings for one beam energy so far
115

    
116
- DC Calibrations?
117
  - Read reports but haven't touched the new ones yet
118
  - Account issues blocked some progress on the DC
119

    
120
Nathan
121
------
122
- High epsilon run plan preparation
123
  - Looked at some candidate settings
124

    
125
Dave G
126
------
127
- FADC programming modification?
128
  - Not done yet
129
  - Want old style data and mode 10 data when it can't find pedestal
130
    - now it just dumps mode 10, so it's difficult to make a comparison 
131
    - this could be particularly helpful for diagnosing problems with KaonLT
132
      data
133

    
134
Stephen
135
-------
136
- Merged Ali's DBASE updates on GitHub
137
- Will merge Jacob's hodoscope updates as soon as they're ready
138
- Diagnose hodoscope batch issues for Jacob
139

    
140
- will do some cdaq checks/updates at the end of the month
141

    
142
- Moved online analysis files from cdaq to the ifarm, all on /cache
143
  - /cache/hallc/c-pionlt/analysis/pionLT-Online_Analysis_Files-2021/
144
    - this should make file access easier than from cdaq
145
    - Will be on tape silo eventually, all on cache now for now
146
    - stubs at /mss/hallc/c-pionlt/...
147

    
148
Next Meeting
149
------------
150
- GH will circulate a poll to select a time for the week of April 18-22
(39-39/412)